OpenStudy (anonymous):

Algera 1 math help please? Terrell and Mindi are organizing a neighborhood skateboarding competition and need to build a mini ramp. Terrell can build the ramp in 12 hours. Mindi can build the ramp in 6 hours. Which shows the equation and solution for the time it will it take Terrell and Mindi to build the ramp if they work together? Round your answer to the nearest hour, if necessary. A) 1 = (1/12 + 1/6)t ; 4 hours B) 1 = (12/1 + 6/1)t ; 3 hours C) 2 = (12/1 + 6/1)t ; 4 hours D) 2 = (1/12 + 1/6)t ; 3 hours

OpenStudy (ash2326):

Terrell can build the ramp in 12 hours, so in 1 hour he will do \(\frac{1}{12}\) of work similarly Mindi will do \(\frac{1}{6}\) of the work in 1 hour, \[\text{So in 1 hour both will do}(\frac{1}{12}+\frac{1}{6} )\text{of the total work}\] \[\text{Therefore total time to build the ramp} =\frac{1}{(\frac{1}{12}+\frac{1}{6} )}\text{hours}\] Can you find this?

OpenStudy (anonymous):

1 = (1/12 + 1/6)t ; 4 hours?

OpenStudy (ash2326):

Yeah